Kelly has a collection of nickels, dimes, and quarters with a total value of $7.30. There are five more nickels than dimes and twice as many quarters as nickels. How many of each coin is in her collection?
? nickels
? dimes
? quarters

Answer by mananth(16946)

dimes x
nickels x+5
quarters2(x+5)
Value 7.30
10x+5(x+5)+25*2(x+5)=730
10x+5x+25+50x+250=730
65x+275=730
65x=730-275
65x=455
/65
x=7 Dimes
Nickels 12
quarters=24